Find the polynomial of lowest degree with the given zeros:
1, 2. I am not understanding how to do this, Please, help. Thanks!

Answer by Edwin McCravy(20054)

Find the polynomial of lowest degree with the given zeros:
1, 2. I am not understanding how to do this, Please, help. Thanks!



Work the problem backwards.

You would end up with

x = 1, x = 2

Before that you would have had:

x - 1 = 0 and x - 2 = 0

Before that you would have had

(x - 1)(x - 2) = 0

And before that you would have had

x² - 3x + 2 = 0

So the simplest polynomial of 
lowest degree you could have had 
would have been:

x² - 3x + 2

Edwin